17. Suppose y varies inversely with x, and y=15 when x=40. Find an equation of variation for x and y.
I appreciate your help in this very much.

Answer by stanbon(75887)

Suppose y varies inversely with x
y = k/x
----------------------------------------
Solve for "k" using "y=15 when x=40"
15 = k/40
k = 600
-----------
Find an equation of variation for x and y.
Equation:
y = 600/x
-----------
Cheers,
Stan H.
--------